# Symmetric Perceptrons, Number Partitioning and Lattices

### v1

@@ Abstract
Our hardness results are versatile, and hold with respect to different distributions of the matrix $\mathbf{A}$ (e.g., i.i.d. uniform entries from $[0,1]$) and weaker requirements on the solution vector $\mathbf{x}$.
@@ 1 Introduction
The problem is versatile, and can be defined with respect to different distributions of the matrix $\mathbf{A}$ (i.i.d. uniform $[0,1]$ entries is another popular choice) and different requirements on the solution vector $\mathbf{x}$. The problem also has rich connections to several other fields including the classical subset sum problem and its variants and the problem of discrepancy minimization.
@@
This comes from the breakthrough work of Bansal [Ban10] and Bansal and Spencer [BS20] from the closely related field of combinatiorial discrepancy theory. [^3] Their result is established for the case of matrices $\mathbf{A}$ with i.i.d. Rademacher entries. Nevertheless, [GKPX22] conjecture that the same guarantee remains true for the case of i.i.d. standard normal entries. Gamarnik, Kızıldağ, Perkins and Xu [GKPX22, GKPX23] studied the large statistical-computational gap scenario in detail and conjectured that no polynomial-time algorithms can achieve a guarantee much better than $\kappa_{\mathsf{comp}}(x)$. In particular, they show that the so-called overlap gap property [MMZ05, AR06, Gl16], which rules out a class of stable algorithms, sets in at
@@
Lattice problems such as the shortest vector problem and the shortest independent vectors problem have been extensively studied for decades largely for their implications to combinatorial optimization [LLL82, Kan83, Kan87, RR23] and even more so, to cryptography [Ajt96, MR07, Reg09]. Time-approximation tradeoffs for lattice problems are well-known [Sch87]. The best known algorithms for these lattice problems employ lattice reduction techniques, based on the LLL and BKZ algorithms [LLL82, Sch87]. They currently all have the following time-approximation trade-off: For a tunable parameter $k$, in dimension $n$, it is posible to solve lattice problems with approximation factor $\gamma=2^{\widetilde{O}(n/k)}$ in time $2^{\widetilde{O}(k)}$, where $\widetilde{O}(\cdot)$ hides polylog factors in $n$. Equalizing these terms gives a $2^{\widetilde{O}(\sqrt{n})}$-time algorithm to solve $2^{\widetilde{O}(\sqrt{n})}$-approximate lattice problems. Despite extensive study in the lattice literature, no better algorithms are known (that improve the exponent by more than a polylog factor). This motivates the assumptions in our theorem statement, both the conservative one and the near-optimal one. (For more discussion on this, see Section 2.2).
Our reduction has several additional features: first, our reduction is versatile and works with respect to different distributions of the matrix $\mathbf{A}$ (e.g., i.i.d. uniform entries from $[0,1]$); and secondly, it shows the hardness not just of computing an SBP solution with $1-o(1)$ probability, but indeed with any non-trivial inverse polynomial probability. For more implications of our reduction, we refer the reader to Section 3.2. Moreover, our full reduction is conceptually simple and direct. We discuss the possibility of using other intermediate problems to establish these same results in Section 1.1.3.
@@
More generally, we can ask the converse questions of the ones raised in our paper. Can lattice algorithms be used to improve algorithms for $\mathrm{NPP}$ or $\mathrm{ABP}$?
@@ 1.1.3 Alternate Reduction Paths
Instead of starting from worst-case lattice problems, we could start with a noisy version of the short integer solutions (SIS) problem in the $\ell_{\infty}$ norm, defined roughly as follows: Given random $\mathbf{A}\sim(\mathbb{Z}/q\mathbb{Z})^{n\times m}$, output $\mathbf{x}\in\mathbb{Z}^{m}\setminus\{\mathbf{0}\}$ such that $\left\|{\mathbf{x}}\right\|_{\infty}$ and $\left\|{\mathbf{A}\mathbf{x}\pmod{q}}\right\|_{\infty}$ are small. The original worst-case to average-case reductions for lattice problems indeed reduce worst-case lattice problems to (average-case) SIS [Ajt96, MR07]. One can reduce (noisy) SIS to $\mathrm{SBP}$ by adding small noise $\mathbf{E}\sim U\left([0,1/q)^{n\times m}\right)$ and setting
@@
(For simplicity, we assume here that $\mathrm{SBP}$ allows matrices with i.i.d. $U([0,1))$ entries instead of standard normal, but we can remove this assumption by sampling from the appropriate discrete Gaussian and scaling appropriately.) Feeding $\mathbf{A}^{\prime}$ into the $\mathrm{SBP}$ solver yields $\mathbf{x}\in\{\pm 1\}^{m}$ such that $\left\|{\mathbf{A}\mathbf{x}}\right\|_{\infty}$ is small, assuming $q\left\|{\mathbf{E}\mathbf{x}}\right\|_{\infty}$ is also sufficiently small.
@@
Brakerski and Vaikuntanathan [BV15] define a problem called the one-dimensional short integer solutions problem (1D-SIS), defined roughly as follows: Given $\mathbf{a}\sim(\mathbb{Z}/q\mathbb{Z})^{m}$, output $\mathbf{x}\in\mathbb{Z}^{m}\setminus\{\mathbf{0}\}$ such that $\left\|{\mathbf{x}}\right\|_{\infty}$ and $|\mathbf{a}^{\top}\mathbf{x}\pmod{q}|$ are small, where $q$ is a product of $n$ primes. [BV15] shows a worst-case to average-case reduction from worst-case lattice problems in dimension $n$ to 1D-SIS. One can reduce 1D-SIS to $\mathrm{NPP}$ by adding small noise $\mathbf{e}\sim U\left([0,1/q)^{m}\right)$ and setting
@@
While these approaches would work, we view this introduction of parameter $q\in\mathbb{Z}$ as extraneous and misleading. To show hardness of a discrete optimization problem in continuous Euclidean space ($\mathrm{SBP}$ and $\mathrm{NPP}$), we should ideally start from a problem that is itself a discrete optimization problem in continuous Euclidean space (worst-case lattice problems). There is no need to add premature discretization by introducing the parameter $q$.
@@ 2 Preliminaries
For a predicate $\varphi$, we use the notation $\mathbf{1}[\varphi]\in\{0,1\}$ to denote the indicator variable of whether $\varphi$ is true ($1$) or false (00). We use $\ln(\cdot)$ to denote the natural logarithm (base $e$) and $\log(\cdot)$ to denote $\log_{2}(\cdot)$. We use $\mathbb{R}_{>0}$ to refer to the set of all positive real numbers, and we use $\mathbb{R}_{\geq 1}$ to refer to the set of all real numbers that are at least $1$. We say a function $f:\mathbb{N}\to\mathbb{R}_{>0}$ is negligible if for all $c\in\mathbb{N}$,
@@ 2.1 Norms & Matrices
We use the notation $\mathbf{0}$ to denote the all 00s vector in $\mathbb{R}^{n}$, where the dimension $n$ is clear from context. We use $\mathbf{I}_{n}\in\mathbb{R}^{n\times n}$ to denote the $n$-dimensional identity matrix. We use the standard $\ell_{1},\ell_{2},\ell_{\infty}$ norms on $\mathbb{R}^{n}$. For a matrix $\mathbf{A}\in\mathbb{R}^{n\times m}$, we use the notation $\sigma_{\max}(\mathbf{A})$ to denote the spectral norm, or maximum singular value, of $\mathbf{A}$. Explicitly,
@@ 2.2 Lattices
For a parameter $\gamma:\mathbb{N}\to\mathbb{R}_{\geq 1}$, the gap covering radius problem ($\mathrm{GapCRP}_{\gamma}$) is a (worst-case) decision problem defined as follows. Given an invertible basis $\mathbf{B}\in\mathbb{R}^{n\times n}$ and threshold $\theta\in\mathbb{R}_{>0}$ as input, output $1$ if $\nu(\mathcal{L}(\mathbf{B}))\leq\theta$ and 00 if $\nu(\mathcal{L}(\mathbf{B}))>\gamma(n)\cdot\theta$.
@@
Definition 4 differs from [MR07, Definition 5.6] in two ways. First, instead of using $\lambda_{n}(\mathcal{L}(\mathbf{B}))$, [MR07] considers general functions $\phi$ mapping $n$-dimensional lattices to $\mathbb{R}_{>0}$. Second, instead of fixing the constant $8$, [MR07] parametrizes this more generally by some constant $g$.
@@
The best known algorithms for these lattice problems employ lattice reduction techniques, based on LLL and BKZ [LLL82, Sch87]. They currently all have the following time-approximation trade-off: For a tunable parameter $k$, in dimension $n$, it is posible to solve lattice problems with approximation factor $\gamma=2^{\widetilde{O}(n/k)}$ in time $2^{\widetilde{O}(k)}$, where $\widetilde{O}(\cdot)$ hides polylog factors in $n$. Equalizing these terms gives a $2^{\widetilde{O}(\sqrt{n})}$-time algorithm to solve $2^{\widetilde{O}(\sqrt{n})}$-approximate lattice problems. Despite extensive study in the lattice literature, no better algorithms are known (that improve the exponent by more than a polylog factor). In particular, the following two assumptions stand:
@@ 2.3 Continuous and Discrete Gaussian Measures
where the last inclusion comes from the bound $\epsilon<1/2$. Subtracting by $1$ and invoking Lemma 6 with $\Lambda=\mathbb{Z}^{n}$ gives the desired bound. ∎
@@ 2.5 Worst-case to Average-case Reductions
where $r$ denotes the internal randomness of $M$. We emphasize that the polynomial in the $\mathsf{poly}(T)$ run-time of $M^{(\cdot)}$ may depend on the non-negligible function $\mu$. Since $x$ is worst-case and $A\in\mathsf{FNP}$, standard amplification applies to make the success probability of $M$ exponentially close to $1$.
@@ 3.1 Proof of Theorem 3
$m$ $=\left\lceil{\left(8\sigma_{2}n^{3/2+\varepsilon}\right)^{1/\varepsilon}}\right\rceil=n^{1+\Theta(1)},$
@@ 3.2 Variants and Generalizations
    Zero entries in $\mathbf{x}$. Instead of requiring $\mathbf{x}\in\{\pm 1\}^{m}$ from the $\mathrm{SBP}$ solver, if we instead allowed $\mathbf{x}\in\{-1,0,1\}^{m}$ with $\mathbf{x}\neq\mathbf{0}$ from the $\mathrm{SBP}$ solver, a similar reduction to the one in Theorem 3 would work as well. The main difference is that the reduction would instead “guess” a coordinate $j\in[m]$ for which $x_{j}\neq 0$ (with success probability at least $1/m$) and put the vector $\mathbf{t}$ in the mean of that coordinate, instead of the first coordinate. (See [MR07] for more rigorous details.)
@@ 4 Reduction to Number Partitioning
For all sufficiently large $N\in\mathbb{N}$, there exist at least $N/\ln(N)$ distinct prime numbers in the interval $[N,10N]$.
@@
for sufficiently large $N$. Therefore, for sufficiently large $N$, by taking the difference of the two quantities, there are at least $N/\ln N$ primes in $[N,10N]$. ∎
@@
Let $p_{1},\dots,p_{n}$ be $n$ distinct prime numbers in the range $[32nm,320nm]$, which we know must exist for sufficiently large $n,m$ by Lemma 14 (since $m\geq n$). Let $q=\prod_{i=1}^{n}p_{i}\leq(320nm)^{n}$. Sample $\mathbf{u}_{1}\sim\operatorname{\mathcal{N}}(\mathbf{t},\sigma_{1}^{2}\mathbf{I}_{n})$, $\mathbf{u}_{2},\cdots,\mathbf{u}_{m}\sim_{\mathrm{i.i.d.}}\operatorname{\mathcal{N}}(\mathbf{0},\sigma_{1}^{2}\mathbf{I}_{n})$. Let $\mathbf{U}=[\mathbf{u}_{1},\mathbf{u}_{2},\cdots,\mathbf{u}_{m}]\in\mathbb{R}^{n\times m}$. Sample $m$ uniformly random lattice vectors $\mathbf{v}_{i}\in\mathcal{L}(\mathbf{B})\mod\mathcal{P}(\mathbf{S})$ (see [Mic04, Proposition 2.9]), and let $\mathbf{V}=[\mathbf{v}_{1},\mathbf{v}_{2},\cdots,\mathbf{v}_{m}]\in\mathbb{R}^{n\times m}$. Define
@@
It suffices to upper bound $\left\|{\widetilde{\varphi}^{-1}(e^{\prime\prime})}\right\|_{1}$ by $1/16$. Recall from Lemma 12 that

### v2

@@ Abstract
Our hardness results are versatile, and hold with respect to different distributions of the matrix $\mathbf{A}$ (e.g., i.i.d. uniform entries from [0,1][0,1][ 0 , 1 ]) and weaker requirements on the solution vector $\mathbf{x}$.
@@ 1 Introduction
The problem is versatile, and can be defined with respect to different distributions of the matrix $\mathbf{A}$ (i.i.d. uniform [0,1][0,1][ 0 , 1 ] entries is another popular choice) and different requirements on the solution vector $\mathbf{x}$. The problem also has rich connections to several other fields including the classical subset sum problem and its variants and the problem of discrepancy minimization.
@@
This comes from the breakthrough work of Bansal [Ban10] and Bansal and Spencer [BS20] from the closely related field of combinatorial discrepancy theory. [^3] Their result is established for the case of matrices $\mathbf{A}$ with i.i.d. Rademacher entries. Nevertheless, [GKPX22] conjecture that the same guarantee remains true for the case of i.i.d. standard normal entries. Gamarnik, Kızıldağ, Perkins and Xu [GKPX22, GKPX23] studied the large statistical-computational gap scenario in detail and conjectured that no polynomial-time algorithms can achieve a guarantee much better than $\kappa_{\mathsf{comp}}(x)$. In particular, they show that the so-called overlap gap property [MMZ05, AR06, Gl16], which rules out a class of stable algorithms, sets in at
@@
Lattice problems such as the shortest vector problem and the shortest independent vectors problem have been extensively studied for decades largely for their implications to combinatorial optimization [LLL82, Kan83, Kan87, RR23] and even more so, to cryptography [Ajt96, MR07, Reg09]. Time-approximation tradeoffs for lattice problems are well-known [Sch87]. The best known algorithms for these lattice problems employ lattice reduction techniques, based on the LLL and BKZ algorithms [LLL82, Sch87]. They currently all have the following time-approximation trade-off: For a tunable parameter $k$, in dimension $n$, it is possible to solve lattice problems with approximation factor $\gamma=2^{\widetilde{O}(n/k)}$ in time $2^{\widetilde{O}(k)}$, where $\widetilde{O}(\cdot)$ hides polylog factors in $n$. Equalizing these terms gives a $2^{\widetilde{O}(\sqrt{n})}$-time algorithm to solve $2^{\widetilde{O}(\sqrt{n})}$-approximate lattice problems. Despite extensive study in the lattice literature, no better algorithms are known (that improve the exponent by more than a polylog factor). This motivates the assumptions in our theorem statement, both the conservative one and the near-optimal one. (For more discussion on this, see Section 2.2).
Our reduction has several additional features: first, our reduction is versatile and works with respect to different distributions of the matrix $\mathbf{A}$ (e.g., i.i.d. uniform entries from [0,1][0,1][ 0 , 1 ]); and secondly, it shows the hardness not just of computing an SBP solution with $1-o(1)$ probability, but indeed with any non-trivial inverse polynomial probability. For more implications of our reduction, we refer the reader to Section 3.2. Moreover, our full reduction is conceptually simple and direct. We discuss the possibility of using other intermediate problems to establish these same results in Section 1.1.3.
@@
For $\mathrm{SBP}$, one can ask what happens in the setting where $m=\Theta(n)$, i.e., so the ratio $m/n$ is $\Theta(1)$. The proof of our hardness result crucially needs $m\geq n^{\Theta(1/\epsilon)}$, with no setting of our parameters yielding $m=\Theta(n)$. Can one extend hardness like in Theorem 1 to this proportional setting?
@@
More generally, we can ask the converse questions of the ones raised in our paper. Can lattice algorithms be used to improve algorithms for $\mathrm{NPP}$, $\mathrm{SBP}$ or $\mathrm{ABP}$? Recent work has given some hope along these lines, showing that lattice-based methods outperform sum-of-squares [ZSWB22].
@@
Related work.
Recently, various optimization and learning problems have been shown to have lattice-based hardness. The formative work of Bruna, Regev, Song and Tang [BRST21] defined a variant of the standard cryptographic Learning With Errors (LWE) problem called Continuous Learning With Errors (CLWE), which is an average-case, periodic, linear regression task, which they showed is as hard as worst-case lattice problems. This and subsequent works have used CLWE as an intermediate problem to show worst-case lattice hardness of learning mixtures of Gaussians, learning single periodic neurons, learning halfspaces, detecting planted backdoors in certain machine learning models, and more [BRST21, SZB21, GVV22, DKMR22, DKR23, Tie23, GKVZ22].
@@ 1.1.3 Alternate Reduction Paths
Instead of starting from worst-case lattice problems, we could start with a noisy version of the short integer solutions (SIS) problem in the $\ell_{\infty}$ norm, defined roughly as follows: Given random $\mathbf{A}\sim(\mathbb{Z}/q\mathbb{Z})^{n\times m}$, output $\mathbf{x}\in\mathbb{Z}^{m}\setminus\{\mathbf{0}\}$ such that $\left\|{\mathbf{x}}\right\|_{\infty}$ and $\left\|{\mathbf{A}\mathbf{x}\pmod{q}}\right\|_{\infty}$ are small. [^5] One can remove the $\pmod{q}$ constraint to reduce from SIS over $\mathbb{Z}$, and this would work as well. The original worst-case to average-case reductions for lattice problems indeed reduce worst-case lattice problems to (average-case) SIS [Ajt96, MR07]. One can reduce (noisy) SIS to $\mathrm{SBP}$ by adding small noise $\mathbf{E}\sim U\left([0,1/q)^{n\times m}\right)$ and setting
@@
(For simplicity, we assume here that $\mathrm{SBP}$ allows matrices with i.i.d. $U([0,1))$ entries instead of standard normal, but we can remove this assumption by sampling from the appropriate discrete Gaussian and scaling.) Feeding $\mathbf{A}^{\prime}$ into the $\mathrm{SBP}$ solver yields $\mathbf{x}\in\{\pm 1\}^{m}$ such that $\left\|{\mathbf{A}\mathbf{x}}\right\|_{\infty}$ is small, assuming $q\left\|{\mathbf{E}\mathbf{x}}\right\|_{\infty}$ is also sufficiently small.
Another approach for $\mathrm{SBP}$ is to reduce directly from Continuous Learning With Errors (CLWE) [BRST21]. This average-case problem, which is known to be as hard as worst-case approximate lattice problems [BRST21, GVV22], asks to distinguish $(\mathbf{A},\mathbf{s}^{\top}\mathbf{A}+\mathbf{e}^{\top}\pmod{1})$ from $(\mathbf{A},\mathbf{b}^{\top})$, where $\mathbf{A}\sim\operatorname{\mathcal{N}}(0,1)^{n\times m}$, $\mathbf{s}$ is drawn spherically from sufficiently large sphere, $\mathbf{e}$ is drawn from a Gaussian with small variance, and $\mathbf{b}$ is a uniformly random vector mod 111. By using integrality of $\mathbf{x}$ and simultaneous smallness of $\mathbf{x}$ and $\mathbf{A}\mathbf{x}$, one can simply multiply the second argument on the right by $\mathbf{x}$ to distinguish from the uniform distribution modulo 111. However, we find this CLWE approach somewhat unsatisfactory since it would show hardness of $\mathrm{SBP}$ assuming at least one of (a) quantum worst-case hardness of search lattice problems or (b) classical worst-case hardness of decisional lattice problems [BRST21, GVV22]. On the other hand, our main result shows hardness of $\mathrm{SBP}$ under only classical worst-case hardness of search lattice problems. As search lattice problems are known to be harder than decisional ones, our main approach in this work yields a stronger result than the CLWE approach. Indeed, for those familiar with lattice-based cryptography, this should not come as a surprise: the difference between $\mathrm{SBP}$ and CLWE is reminiscent of the difference between SIS and the Learning With Errors (LWE) problems in lattice-based cryptography.
@@
Brakerski and Vaikuntanathan [BV15] define a problem called the one-dimensional short integer solutions problem (1D-SIS), defined roughly as follows: Given $\mathbf{a}\sim(\mathbb{Z}/q\mathbb{Z})^{m}$, output $\mathbf{x}\in\mathbb{Z}^{m}\setminus\{\mathbf{0}\}$ such that $\left\|{\mathbf{x}}\right\|_{\infty}$ and $|\mathbf{a}^{\top}\mathbf{x}\pmod{q}|$ are small, where $q$ is a product of $n$ primes. [^6] Similarly here, one can remove the $\pmod{q}$ constraint and reduce from 1D-SIS over $\mathbb{Z}$. [BV15] shows a worst-case to average-case reduction from worst-case lattice problems in dimension $n$ to 1D-SIS. One can reduce 1D-SIS to $\mathrm{NPP}$ by adding small noise $\mathbf{e}\sim U\left([0,1/q)^{m}\right)$ and setting
@@
While these approaches would work, we view this introduction of parameter $q\in\mathbb{Z}$ or modulus reduction as extraneous and misleading. To show hardness of a discrete optimization problem in continuous Euclidean space ($\mathrm{SBP}$ and $\mathrm{NPP}$), we should ideally start from a problem that is itself a discrete optimization problem in continuous Euclidean space (worst-case lattice problems). There is no need to add premature discretization by introducing the parameter $q$ or unnecessary modular reduction.
@@ 2 Preliminaries
For a predicate $\varphi$, we use the notation $\mathbf{1}[\varphi]\in\{0,1\}$ to denote the indicator variable of whether $\varphi$ is true (111) or false (0). We use $\ln(\cdot)$ to denote the natural logarithm (base $e$) and $\log(\cdot)$ to denote $\log_{2}(\cdot)$. We use $\mathbb{R}_{>0}$ to refer to the set of all positive real numbers, and we use $\mathbb{R}_{\geq 1}$ to refer to the set of all real numbers that are at least 111. We say a function $f:\mathbb{N}\to\mathbb{R}_{>0}$ is negligible if for all $c\in\mathbb{N}$,
@@ 2.1 Norms & Matrices
We use the notation $\mathbf{0}$ to denote the all 0s vector in $\mathbb{R}^{n}$, where the dimension $n$ is clear from context. We use $\mathbf{I}_{n}\in\mathbb{R}^{n\times n}$ to denote the $n$-dimensional identity matrix. We use the standard $\ell_{1},\ell_{2},\ell_{\infty}$ norms on $\mathbb{R}^{n}$. For a matrix $\mathbf{A}\in\mathbb{R}^{n\times m}$, we use the notation $\sigma_{\max}(\mathbf{A})$ to denote the spectral norm, or maximum singular value, of $\mathbf{A}$. Explicitly,
@@ 2.2 Lattices
For a parameter $\gamma:\mathbb{N}\to\mathbb{R}_{\geq 1}$, the gap covering radius problem ($\mathrm{GapCRP}_{\gamma}$) is a (worst-case) decision problem defined as follows. Given an invertible basis $\mathbf{B}\in\mathbb{R}^{n\times n}$ and threshold $\theta\in\mathbb{R}_{>0}$ as input, output 111 if $\nu(\mathcal{L}(\mathbf{B}))\leq\theta$ and 0 if $\nu(\mathcal{L}(\mathbf{B}))>\gamma(n)\cdot\theta$.
@@
Definition 4 differs from [MR07, Definition 5.6] in two ways. First, instead of using $\lambda_{n}(\mathcal{L}(\mathbf{B}))$, [MR07] considers general functions $\phi$ mapping $n$-dimensional lattices to $\mathbb{R}_{>0}$. Second, instead of fixing the constant 888, [MR07] parametrizes this more generally by some constant $g$.
@@
The best known algorithms for these lattice problems employ lattice reduction techniques, based on LLL and BKZ [LLL82, Sch87]. They currently all have the following time-approximation trade-off: For a tunable parameter $k$, in dimension $n$, it is possible to solve lattice problems with approximation factor $\gamma=2^{\widetilde{O}(n/k)}$ in time $2^{\widetilde{O}(k)}$, where $\widetilde{O}(\cdot)$ hides polylog factors in $n$. Equalizing these terms gives a $2^{\widetilde{O}(\sqrt{n})}$-time algorithm to solve $2^{\widetilde{O}(\sqrt{n})}$-approximate lattice problems. Despite extensive study in the lattice literature, no better algorithms are known (that improve the exponent by more than a polylog factor). In particular, the following two assumptions stand:
@@ 2.3 Continuous and Discrete Gaussian Measures
where the last inclusion comes from the bound $\epsilon<1/2$. Subtracting by 111 and invoking Lemma 6 with $\Lambda=\mathbb{Z}^{n}$ gives the desired bound. ∎
@@ 2.5 Worst-case to Average-case Reductions
where $r$ denotes the internal randomness of $M$. We emphasize that the polynomial in the $\mathsf{poly}(T)$ run-time of $M^{(\cdot)}$ may depend on the non-negligible function $\mu$. Since $x$ is worst-case and $A\in\mathsf{FNP}$, standard amplification applies to make the success probability of $M$ exponentially close to 111.
@@ 3.1 Proof of Theorem 3
$m$ $=\left\lceil{\left(8\sigma_{2}n^{3/2+\varepsilon}\right)^{1/\varepsilon}}\right\rceil=n^{\Theta(1/\varepsilon)},$
@@ 3.2 Variants and Generalizations
[…]
@@ 4 Reduction to Number Partitioning
[…]
@@
[…]
@@
[…]
@@
It suffices to upper bound $\left\|{\widetilde{\varphi}^{-1}(e^{\prime\prime})}\right\|_{1}$ by 1/161/161 / 16. Recall from Lemma 12 that